OpenStudy (omgthatgirl):

Rewrite the equation 2y^2+8y-2x+6=0 in standard form A. (y+2)^2=2(x+1) B. (y+4)^2=x+1 C. (y+2)^2=x+1 D. (y+2)^2=x-3

OpenStudy (anonymous):

2y^2 +8y = 2x - 6

OpenStudy (anonymous):

we can divide this by 2

OpenStudy (anonymous):

it will be: y^2 + 4 = x - 3

OpenStudy (omgthatgirl):

Ok

OpenStudy (anonymous):

what u will do next?

OpenStudy (omgthatgirl):

(Y+2)^2?

OpenStudy (anonymous):

+

OpenStudy (omgthatgirl):

3

OpenStudy (anonymous):

??

OpenStudy (omgthatgirl):

4?

OpenStudy (anonymous):

D?

OpenStudy (omgthatgirl):

Oh ok

OpenStudy (anonymous):

i ask u